# DeepEdit: Knowledge Editing as Decoding with Constraints

### v4

@@ Abstract
How to edit the knowledge in multi-step reasoning has become the major challenge in the knowledge editing (KE) of large language models (LLMs). The difficulty arises because the hallucinations of LLMs during multi-step reasoning often lead to incorrect use of new knowledge and incorrect answers. To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge. We propose a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs’s ability to generate coherent reasoning chains with new knowledge through depth-first search. Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce succinct and coherent reasoning chains involving new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@ 1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to incorporate new or updated knowledge, in contrast to relying solely on outdated parametric knowledge. Multi-hop question answering is the challenging task to evaluate KE Zhong et al. (2023), which requires models to answer questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
Figure 1: (a) An instance of multi-hop question answering with new knowledge. (b) Prior KE methods, e.g., MeLLo Zhong et al. (2023), suffers from the hallucinations of LLM generators. (c) Our DeepEdit controls LLMs with decoding constraints to produce coherent reasoning chains involving new knowledge.
@@
Recent literature has mainly focused on performing KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023), which puts new knowledge in the input context to fulfill KE (see Figure 1(b)). What new knowledge to use and how to utilize new knowledge is decided by the LLM generators’ implicit reasoning. However, the LLM generator’s hallucinations against the in-context new knowledge can harm the logical coherence of multi-step reasoning and result in incorrect answers Zhang et al. (2023).
To get the correct answers to multi-hop questions with new knowledge, we should ensure that the adjacent reasoning steps are logically coherent, and new knowledge is properly integrated into reasoning (see Fig. Figure 1(a). In this work, we explore using decoding constraints and search algorithms to explicitly control LLMs’ reasoning with new knowledge.
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints that facilitate LLMs to soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted parametric knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2). A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
We incorporate the above constraints into a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing). When producing every reasoning step, DeepEdit searches through the parametric and new knowledge to find the knowledge items that satisfy our constraints, termed as valid step candidates. Among the valid candidates, DeepEdit prioritizes more important ones as the new reasoning step for further reasoning. We comprehensively evaluate the importance of different candidates based on their semantic relationships and the prior that the new knowledge is more informative to LLMs than the parametric knowledge.
Note that the classical depth-first search has a backtracking mechanism that enables us to traverse all the reasoning chains of valid reasoning steps Tarjan (1972). However, such backtracking is undoubtedly time-consuming and unnecessary after an answer is found. Therefore, we design an early-stopping mechanism to stop the depth-first search when an answer is found by LLMs. Such an early stopping significantly improves the reasoning efficiency by reducing redundant iterations, as validated by the qualitative analysis (see Section 3.3) and empirical results (see Section 5.4).
The advantages of DeepEdit are two-fold. First, DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. Second, DeepEdit improves both the receptiveness of new knowledge and the coherence of multi-step reasoning by directly controlling the LLMs’ reasoning with decoding constraints. Unlike previous KE methods that edit either the model parameters Zhu et al. (2020) or the input prompts Cohen et al. (2023), we directly control the decoding of LLMs to incorporate the new knowledge into outputs, which opens up a new direction for KE in LLMs.
@@ 2 Related Work
Different from the prior work, we propose a new KE method DeepEdit, which effectively augments LLMs on KE with constrained decoding. We illustrate a comparison between DeepEdit and existing KE methods in Table 1. Our DeepEdit outperforms the prior KE methods by editing the LLMs’ outputs directly, which paves the way to a new direction of LLMs’ KE.
@@
Figure 3: When producing every reasoning step, we retrieve the edited facts that are close to the generated step in the semantic space (left). Only the edited facts that satisfy all the constraints, i.e., valid edited facts, are considered as the step candidates (right).
KE Method Editing Editing Editing Model Parameters Input Prompts Outputs MEND Mitchell et al. (2021) MEMIT Meng et al. (2022c) ICE Cohen et al. (2023) MeLLo Zhong et al. (2023) DeepEdit (Ours)
Table 1: Comparison between KE methods.
@@ 3.1 Constraint Designing for Knowledge Editing
    Conciseness: Every step is identical.
@@ 3.2 Constraint Verification on Every Step
At every iteration, we take both the parametric and new knowledge as the step candidates. First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be most useful to answer the target question. Second, we retrieve $N$ new knowledge closest to the generated reasoning step. The distance between knowledge and the generated step is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al. (2019). The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question. We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
We verify the constraints on the aforementioned step candidates to identify the valid ones. In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences. For example, Coherence’s verification can be completed by assessing whether a pair of reasoning steps is relevant to each other or not. Therefore, one option to do constraint verification is to train a binary classifier for every constraint. However, such a solution is not data or time-efficient. In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers. Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from MQuAKE Zhong et al. (2023). In our experiments, we set $D=2$ and $N=2$ by default.
@@ 3.3 Efficient Reasoning with Depth-first Search
At every iteration, there may exist more than valid step candidates. We can consider each valid step candidate simultaneously as the new reasoning step and make further reasoning based on each of them separately. This design can be seen as a breadth-first search of a valid multi-hop reasoning chain. It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps. This raises a natural question:
@@ 4 New Benchmarks for More Precise and Challenging Evaluation of KE
We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard. Both are built based on MQuAKE Zhong et al. (2023), a recent KE dataset that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge. Every instance in MQuAKE includes a multi-hop question and several edited facts, every of which can alter the ground-truth answer to the question.
@@ 5.1 Experimental Settings
We take the following KE benchmarks MQuAKE-3k Zhong et al. (2023), MQuAKE-2002, and MQuAKE-hard for evaluation. We follow the existing work Zhong et al. (2023) to use accuracy as the main evaluation metric. For every instance, we test the accuracy of the first question in the instance, since different questions in the same instance are semantically equivalent and correspond to the same ground-truth answer. We follow Zhong et al. (2023) to set the KE batch size as 100 and full batch for evaluation. The KE batch size means the number of instances that provide new knowledge for retrieval when answering every question.
@@ 6 Conclusion
Acknowledgements
This work was partially supported by a DARPA ANSR program FA8750-23-2-0004. The views and conclusions are those of the authors and should not reflect the official policy or position of DARPA or the U.S. Government.
@@ Appendix C Case Study
Figure 10: Text Generation Prompt of DeepEdit Zhong et al. (2023)

### v5

@@ Abstract
Incorporating new knowledge in multi-step reasoning has become a major problem in the research of knowledge editing (KE). LLMs often struggle to generate coherent and precise reasoning chains when incorporating new in-context knowledge, as they tend to deviate from relevant information and lose logical coherence during reasoning. To address this issue, it is crucial to apply direct controls over LLM outputs to ensure both logical coherence and effective integration of new knowledge into their reasoning processes. To this end, we propose DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing), a KE method designed to enhance LLMs’ reasoning capabilities with new knowledge. DeepEdit employs the decoding constraints that we carefully crafted to explicitly improve logical coherence and knowledge integration during reasoning. Additionally, we introduce a semantic-based importance estimator that ranks the relevance of reasoning steps to optimize the informativeness of output reasoning relative to the input question. Beyond DeepEdit, we propose two new KE benchmarks, MQuAKE-2002 and MQuAKE-hard, which offer more precise and challenging assessments of KE methodologies. Qualitative analysis shows that DeepEdit enables LLMs to generate succinct and coherent reasoning chains incorporating new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@ 1 Introduction
The research of knowledge editing (KE) aims to enable large language models (LLMs) to be such question-answering systems that the outputs are receptive to the new knowledge that is different from the LLMs’ parametric knowledge. Multi-hop question answering with new knowledge is currently the major task to evaluate KE Zhong et al. (2023), which requires models to answer questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
Figure 1: (a) An instance in KE evaluation benchmarks. (b,c) Prior KE methods that retrain LLMs with new knowledge (model editing) and adds new knowledge in the input prompt (in-context editing). (c) Our DeepEdit directly controls LLMs’ outputs to select the valid and most important (new) knowledge for reasoning.
Existing literature has mainly focused on performing KE through model editing Mitchell et al. (2021); Meng et al. (2022b) and in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023). The former uses retraining to inject new knowledge into LLMs’ weights while the latter adds new knowledge to the input prompt to fulfill KE (see Figure 1(b)). Both of them lack direct control of LLMs’ outputs, which is essential to enhance LLMs’ multi-step reasoning incorporating new knowledge (see Figure 2).
A satisfactory multi-step reasoning chain with new knowledge should consist of coherent steps and place new knowledge at appropriate positions (see Fig. Figure 1(a). In this work, we explore to propose a new constrained decoding algorithm that enhances LLMs’ to produce such reasoning chains. First, we design the decoding constraints that facilitate LLMs to soundly incorporate new knowledge in multi-step reasoning. Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted parametric knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 3).
Reasoning without New Knowledge Question: The writer of the novel “1984” died in which continent? Reasoning Steps: The writer of the novel "1984" is George Orwell. George Orwell died in the city of London. London is in the continent of Europe. Answer: Europe ✓
Reasoning with New Knowledge in Context Question: Assume George Orwell died in the city of Bucharest; The writer of “Harry Potter” is Stephen King; Bucharest is located in the continent of North America; Berlin is located in the continent of Asia. The writer of the novel “1984” died in which continent? Reasoning Steps: The writer of the novel "1984" is George Orwell. The writer of “Harry Potter” is Stephen King. George Orwell died in the city of Bucharest. Berlin is located in the continent of Asia. Answer: Asia ✗
Reasoning with New Knowledge in Decoding (Ours) Question: The writer of the novel “1984” died in which continent? Reasoning Steps: The writer of the novel "1984" is George Orwell. George Orwell died in the city of Bucharest. Bucharest is located in the continent of North America. Answer: North America ✓
Figure 2: (Upper) It is easy for LLMs to generate coherent reasoning chains without new knowledge. (Middle) Given new knowledge in context, LLMs can hardly generate coherent reasoning chains due to hallucinations on new knowledge. As a result, both relevant and irrelevant knowledge participate in the reasoning. (Lower) We do not put new knowledge in context but let them directly contribute to the decoding, which leads to coherent and precise reasoning chains.
We incorporate the above constraints into a new KE method: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing). When producing every reasoning step, DeepEdit searches through the parametric and new knowledge to find the knowledge items that satisfy our constraints, termed as valid step candidates. Among the valid candidates, DeepEdit prioritizes more important ones as the new reasoning step for further reasoning. We comprehensively evaluate the importance of different candidates based on their semantic relationships and the prior that the new knowledge is more informative to LLMs than the parametric knowledge.
Classical depth-first search has a backtracking mechanism that enables us to traverse all the reasoning chains of valid reasoning steps Tarjan (1972). Such backtracking is time-consuming and unnecessary after an answer is found. Therefore, we design an early-stopping mechanism to stop the depth-first search when an answer is found by LLMs. Such an early stopping significantly improves the reasoning efficiency by reducing redundant iterations, as validated by the qualitative analysis (see Section 3.3) and empirical results (see Section 5.4).
DeepEdit is flexibly applicable to any black-box LLM that only requires access to output texts. Besides, DeepEdit improves both the receptiveness of new knowledge and the coherence of multi-step reasoning by directly controlling the LLMs’ reasoning with decoding constraints. Unlike previous KE methods that edit either the model parameters Zhu et al. (2020) or the input prompts Cohen et al. (2023), we directly control the output of LLMs to incorporate the new knowledge, which opens up a new direction for KE in LLMs.
@@ 2 Related Work
Different from the prior work, we propose a new KE method DeepEdit, which effectively augments LLMs on KE with constrained decoding. Our DeepEdit outperforms the prior KE methods by editing the LLMs’ outputs directly, which paves the way to a new direction of LLMs’ KE.
@@ 3.1 Constraint Designing for Knowledge Editing
    Conciseness: Every step is unique.
@@ 3.2 Constraint Verification on Every Step
At every iteration, we take both the parametric and new knowledge as the step candidates. First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be most useful to answer the target question. Second, we retrieve $N$ new knowledge closest to the generated reasoning step. The distance between knowledge and the generated step is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al. (2019). The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question. We take these $N+1$ knowledge as the reasoning step candidates.
We verify the constraints on the aforementioned step candidates to identify the valid ones. In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences. For example, Coherence’s verification can be completed by assessing whether a pair of reasoning steps is relevant to each other or not. Therefore, one option to do constraint verification is to train a binary classifier for every constraint. However, such a solution is not data or time-efficient. In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers. Every verifier’s prompt includes $D$ positive and $N$ negative demonstrations randomly sampled from MQuAKE Zhong et al. (2023). In our experiments, we set $D=2$ and $N=2$ by default.
@@ 3.3 Efficient Reasoning with Depth-first Search
At every iteration, there may exist more than one valid step candidates. We can consider each valid step candidate simultaneously as the new reasoning step and make further reasoning based on each of them separately. This design can be seen as a breadth-first search of a valid multi-hop reasoning chain. It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps. This raises a natural question:
@@ 4 New Benchmarks for More Precise and Challenging Evaluation of KE
We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard. Both are built based on MQuAKE Zhong et al. (2023), a recent KE dataset that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge. Every instance in MQuAKE includes a multi-hop question and several edited facts, each of which can alter the ground-truth answer to the question.
@@ 5.1 Experimental Settings
We take the following KE benchmarks MQuAKE-3k Zhong et al. (2023), MQuAKE-2002, and MQuAKE-hard for evaluation. We follow the existing work Zhong et al. (2023) to use accuracy as the main evaluation metric. For every instance, we test the accuracy of the first question in the instance, since different questions in the same instance are semantically equivalent and correspond to the same ground-truth answer. We follow Zhong et al. (2023) to set the KE batch size as 100 and full batch for evaluation. The KE batch size means the number of instances that provide new knowledge for retrieval when answering each question.
@@ 6 Conclusion
Limitations
DeepEdit is an early effort at conducting knowledge editing in the decoding stage. There still exist spaces for improving DeepEdit. For example, can we have more decoding constraints to improve knowledge editing? And can we develop more advanced decoding strategies to make LLMs follow the constraints? We hope the future work can answer above questions and further improve knowledge editing to serve more real-world applications.
Ethical Considerations
Ethical considerations are of utmost importance in our research endeavors. In this paper, we strictly adhere to ethical principles by exclusively utilizing open-source datasets and employing various models that are either open-source or widely recognized in the scientific community. We are committed to upholding ethical standards throughout the research process, prioritizing transparency, and promoting the responsible use of technology for the betterment of society.
@@ Appendix C Case Study
Appendix D Discussion with MeLLo
MeLLo Zhong et al. (2023) utilizes a prompt engineering based greedy search to produce a reasoning chain that is receptive to new knowledge. At every iteration, MeLLo retrieves a new knowledge item that is semantically close to the generated reasoning step. It assesses whether the retrieved new knowledge is conflicted with the generated step. If so, it takes the new knowledge as next reasoning step, or the generated step otherwise. Below, we compare the differences between MeLLo and our DeepEdit from different perspectives:
  • Generation Prompt. As shown in Fig. Figure 10 and Figure 11, MeLLo’s generation prompt is much more complex than our DeepEdit. MeLLo leads to much longer context than our DeepEdit for every inference, which induces more serious hallucinations and time/computational costs.
  • Greedy Search versus Depth-First Search. MeLLo employs the greedy search to find the reasoning chain consisting of both parametric and new knowledge. If there is error on any reasoning step, the final answer cannot be gotten without the backtracking mechanism. We employ the depth-first search to look for satisfactory reasoning chains with new knowledge, of which the backtracking mechanism provides a fault-tolerance mechanism to complement the mistakes happening in any reasoning steps.
  • New Knowledge Selection How and where to use the new knowledge in the reasoning chain is the key to getting the correct answer with new knowledge. MeLLo only uses a conflict assessment defined in the prompt template to determine whether to use a new knowledge item. In contrast, our DeepEdit comprehensively evaluates different new knowledge using explicit constraints and a semantic-based importance estimator.
@@ Appendix D Discussion with MeLLo
Figure 10: Text Generation Prompt of DeepEdit (Ours).